OpenStudy (anonymous):

how do i find the arc cot(10/23) in radians and round to 2 decimals

OpenStudy (anonymous):

\[\frac{1}{\tan^{-1}(10/23)}\]

OpenStudy (anonymous):

thank you so much!

OpenStudy (anonymous):

I would have guessed that my HP 50g graphing calculator had an arccot function key, but it doesn't. With the value of 10/23 showing in the display, your push the 1/x key and then the arctan key. Use Mathematica most of the time on this site and if you type in "ArcCot[10./23]" and hit shift-Enter, 1.16067 is returned to the screen.

OpenStudy (anonymous):

my answer wasn't correct? :( what did i do wrong? i have a ti83plus

OpenStudy (anonymous):

I think that Tyler1992 has it wrong.\[\text{ArcCot }\left[\frac{10}{23}\right]\text{=}\text{ArcTan }\left[\frac{23}{10}\right]\]